OpenStudy (anonymous):

express in the standard form a+bi. (1/6+sqrt19/6i)^2

OpenStudy (anonymous):

you can use \[(a+b)^{2} = a ^{2} + b ^{2} + 2ab\] \[(1/6 +\sqrt{19/6i})^{2} = (1/6)^{2} + (\sqrt{19/6i})^{2} + 2(1/6)(\sqrt{19/6i})\]